# Transformer-based stereo-aware 3D object detection from binocular images

### v3

@@ Abstract
Transformers have shown promising progress in various visual object detection tasks, including monocular 2D/3D detection and surround-view 3D detection. More importantly, the attention mechanism in the Transformer model and the image correspondence in binocular stereo are both similarity-based. However, directly applying existing Transformer-based detectors to binocular stereo 3D object detection leads to slow convergence and significant precision drops. We argue that a key cause of this defect is that existing Transformers ignore the stereo-specific image correspondence information. In this paper, we explore the model design of Transformers in binocular 3D object detection, focusing particularly on extracting and encoding the task-specific image correspondence information. To achieve this goal, we present TS3D, a Transformer-based Stereo-aware 3D object detector. In the TS3D, a Disparity-Aware Positional Encoding (DAPE) module is proposed to embed the image correspondence information into stereo features. The correspondence is encoded as normalized sub-pixel-level disparity and is used in conjunction with sinusoidal 2D positional encoding to provide the 3D location information of the scene. To extract enriched multi-scale stereo features, we propose a Stereo Preserving Feature Pyramid Network (SPFPN). The SPFPN is designed to preserve the correspondence information while fusing intra-scale and aggregating cross-scale stereo features. Our proposed TS3D achieves a 41.29% Moderate Car detection average precision on the KITTI test set and takes 88 ms to detect objects from each binocular image pair. It is competitive with advanced counterparts in terms of both precision and inference speed.
@@ I Introduction
Stereo vision systems can perceive 3D scene for autonomous vehicles and intelligent transportation [1]. In binocular stereo based 3D perception, image correspondence information plays an essential role [2, 3, 4]. Image correspondence information, in the form of disparity or depth supervision, is widely used in convolutional neural network (CNN) based binocular 3D object detection to prevent the binocular detector from converging to a monocular local minimum [5, 6].
Vision Transformers [7], which are equipped with similarity-based attention mechanisms and long-range modeling ability, seem compatible with binocular detection at first glance. On the one hand, the attention mechanism in Transformers and the critical image correspondence information in binocular stereo are both similarity-based. On the other hand, the Transformer’s inherent ability of long-range modeling suites for capturing the wide range of disparities (e.g., a range of $(0,192]$ in KITTI [8]). In addition, Transformers have gained promising progress in 2D object detection [9, 10, 11], surround-view 3D object detection [12, 13, 14], and depth estimation [15]. Therefore, we focus on adopting the Transformer model to binocular 3D detection in this paper, which can hopefully benefit future study by enabling binocular detection to interact with Transformer-based language models. However, for binocular 3D object detection in driving scenes, to the best of our knowledge, there have not been any public Transformer-based detectors in the literature.
@@
An intuitive way to introduce Transformer models into binocular 3D object detection is to adopt the above surround-view Transformer-based 3D detectors to the binocular scene (e.g., KITTI [8]). To achieve that, we train binocular revisions of existing surround-view 3D object detectors, which are named with Binocular suffices. Their validation APs after training 320 epochs are enumerated in Fig. 1b, and the validation AP curve of DETR3D-Binocular during training is shown in Fig. 1a as blue circles. It can be seen that existing Transformer detectors converge to a poor 3D detection performance, which can not be compared to advanced CNN-based binocular 3D object detectors nor to their surround-view detection performances [16].
We argue that the convergence problem is caused by the lack of image correspondence information in surround-view 3D object detection Transformers such as DETR3D, as well as in existing monocular-based 3D object detection Transformers. It is a fact that monocular images do not involve cross-view correspondence [17, 18]; and the overlapping area across images in a surround-view system is small [16]. Concerning the binocular vision system, however, the overlap areas between left and right images occupy a large proportion of the pixels. The latent image correspondence information in the large overlap areas is crucial for 3D scene perception. Therefore, extracting and exploiting the image correspondence information in a Transformer-based model is a task-specific challenge of binocular 3D object detection.
@@
Decoder-wisely, a deformable Transformer decoder [21] is introduced to decode multi-scale stereo features. In the decoder, the DAPE is designed to explicitly encode the sub-pixel-level correspondence information into the stereo features. DAPE composites normalized disparities and sinusoidal 2D positional encoding [21], thus simultaneously providing position information in both 2D image space and 3D scene. In addition to the DAPE, a Non-Parametric Anchor-Based Object Query (NPAQuery) scheme is introduced. Compared with widely-used parametric queries, NPAQuery does not introduce new learnable parameters as query embeddings, instead, it reuses the stereo features as a uniformly distributed query embedding in the 2D image space.
Experiments on the KITTI dataset show that TS3D is an effective Transformer model for binocular 3D object detection, and it is competitive with existing CNN-based counterparts in both terms of accuracy and efficiency. Consequently, our TS3D can encode the image correspondence information in Transformer model, thus alleviating the above convergence problem. The validation AP curve of our TS3D is plotted in Fig. 1a as orange triangles, and its validation results are listed in Fig. 1b. It is shown that after preserving and encoding the image correspondence information by the proposed SPFPN and DAPE respectively, the performance gap is significantly diminished. To the best of our knowledge, the proposed TS3D is the first public Transformer-based binocular 3D object detector for driving scenes in the literature.
@@
    We present the Transformer-based Stereo-aware 3D object detector (TS3D), which is the first public Transformer-based binocular detector in the literature and can preserve and encode the image correspondence information from binocular images.
@@ II Related work
Existing binocular 3D object detectors are based on convolutional neural networks (CNN) and are categorized in this section. Representative Transformer-based 3D detectors for monocular images, LiDAR point clouds, and surround-view images are briefly introduced, and readers are encouraged to refer to [22, 23] for more thorough reviews on Transformers-based 3D object detection. Existing feature pyramid networks are reviewed as well, which relates to our SPFPN.
Existing binocular 3D object detectors can be divided into two categories based on their sources of the image correspondence information: detectors with and without LiDAR-based image correspondence supervision. The former [24, 6, 25], shortened as Stereo-with-LiDAR, generates ground-truth image correspondence information based on projected LiDAR point clouds, resulting in accurate yet sparse supervision; whereas the latter [26, 27, 28], shortened as Stereo-without-LiDAR, does not rely on LiDAR point clouds to generate ground-truth image correspondence information. However, the pseudo ground-truths obtained by external stereo matching networks can provide inaccurate supervision in those methods [5, 29]. Some detectors in this category do not use explicit image correspondence supervision [30, 31]. Although Stereo-with-LiDAR detectors achieve higher detection precision, developing Stereo-without-LiDAR detectors is still essential as LiDAR devices may not always be affordable or available, such as in stereo endoscopes and traffic cameras.
@@
The Transformer-based 2D object detector DETR [9] re-formulates the 2D object detection as a collection prediction task and removes the anchor generation and Non-Maximum Suppression (NMS) processes, resulting in an end-to-end detector. DeformableDETR [21] introduces a deformable attention module, which merely applies the multi-head attention mechanism at several learnable key locations around the reference point. It significantly reduces the computational cost of the DETR decoder and converges faster. The Transformer-based unary 3D object detector MonoDTR [17] utilizes auxiliary supervision to implicitly learn depth-aware features and a depth-aware Transformer to mine context and depth information from global features. MonoDETR [43] takes 3D object candidates as queries and introduces an attention-based depth encoder to encode depth embeddings. A depth cross-attention module is proposed to decode inter-query and query-scene interactions. The Transformer-based surround-view 3D object detector DETR3D [12] uses a CNN to extract unary features of each view and a DETR decoder to predict objects. PETR [13] extends DETR3D by generating a 3D coordinate grid in the camera frustum space and fusing the 3D coordinates with the 2D features. BEVFormer [14] generates object queries on the bird’s eye view (BEV) and uses both spatial and temporal information. A spatial cross-attention module is proposed to extract BEV features from each view and a temporal self-attention module is designed to fuse historical BEV features. In addition to the obvious differences in input modalities comparing our TS3D with the above Transformer-based detectors, the image correspondence information are preserved and encoded into enriched stereo feature pyramids by two novel components, namely SPFPN and DAPE.
@@
Feature Pyramid Network (FPN) [44] exploits the multi-level nature of CNNs to generate multi-scale features, so that the features of each level focus more on objects of the specific scale. FPN improves the detection performance of multi-scale 2D object detection. RetinaNet [45] makes use of more levels of the features than that used in FPN, resulting in higher detection precision at a lower computational cost. Original FPN fuses features from lower-resolution to higher-resolution (i.e., top-down), and PAFPN [46] adds an aggregation network from higher-resolution to lower-resolution features (i.e., bottom-up). ASFF [47] adaptively fuses multi-scale features and spatially filters information that conflicts across scales. BiFPN [48] designs a simple and efficient bidirectional FPN, in which a basic structure consisting of top-down and bottom-up paths is repeated to capture enriched multi-scale features. RFP [49] recursively reuses the FPN structure and introduces Atrous Spatial Pyramid Pooling (ASPP) module as a connection module between two FPNs, which improves the performance of 2D object detection. The above feature pyramids are carefully designed for unary 2D object detection and do not consider the image correspondence nature of binocular images. As a result, they can be applied to generate unary feature pyramids in binocular 3D object detection, nevertheless, directly applying them in stereo features will destroy the latent image correspondence information and affect the accuracy of 3D object detection.
@@ III-A Overall architecture of TS3D
The above unary features are then fed into Stereo Preserving Feature Pyramid Network (SPFPN). Both primary and enhanced unary feature pyramids are exploited to construct correlation-based cost volumes [51], resulting in primary and enhanced cost volume pyramids (shown as orange boxes in the second column of Fig. 2). Constructed cost volume pyramids are then referred to as stereo feature pyramids in the SPFPN. Subsequently, primary and enhanced stereo features from each level of the two pyramids are merged, in which the correspondence information is kept intact. The resultant stereo-preserving feature pyramid is then cross-scale aggregated in a bottom-up manner. Finally, enriched and aggregated multi-scale stereo features are obtained for subsequent disparity estimation and Transformer decoder.
@@
The Transformer decoder consists of Multi-Head Self-Attention layers [9] (shortened as MHSA in Fig. 2), Multi-Scale Deformable Cross-Attention layers [21] (MSDeformCA), and Feed-Forward Networks (FFNs). The last feature tensor for disparity regression is termed disparity logits and is used as the input of our DAPE. The generated positional encoding of the 3D scene is summed with the query of each layer in the Transformer decoder. The lowest-resolution features in the above feature pyramid are convolved to generate embeddings of the NPAQuery. In addition to the DAPE and NPAQuery, the multi-scale stereo features extracted by SPFPN are exploited as keys and values of the MSDeformCA layers. The above processes including MHSA, MSDeformCA, and FFN are collectively referred to as a decoder layer. The Transformer decoder is a cascade of $N_{\text{dec}}$ decoding layers, in which the object queries are refined step-by-step. During training, 3D objects are classified and regressed under ground-truth supervision after each decoder layer; During inference, only predictions from the last layer are taken as the final 3D object detection results.
@@ III-B Stereo Preserving Feature Pyramid Network
Refer to caption (a) FPN Refer to caption (b) BiFPN
@@ III-C1 Disparity-Aware Positional Encoding (DAPE)
It can be seen from Eqs. 5 and 6 that the DAPE not only encodes 2D information in the 2D image space but also explicitly encodes the correspondence information of binocular images. It is suitable for locating the projection of 3D object center points and bounding boxes in 2D images; The encoded image correspondence information is guided by stereo features, thus is suitable for regressing depths and sizes of 3D objects in 3D space.
@@ IV-A Experimental setup
The experiments are conducted on the KITTI dataset [8]. In the ablation experiments, the data split is consistent with that of [55] and the evaluation index is obtained on its validation subset. We report all difficulty levels (Easy, Moderate, Hard) of 3D Car detection AP (Average Precision) as benchmarking indicators. In the experiments in Sec. IV-C, the entire training set is used during training, and the prediction results are calculated by the KITTI server. Detection APs of three difficulty levels of 3D Car and Pedestrian detection tasks are reported on KITTI test set. Among them, Moderate 3D Car detection AP is the main evaluation index in this paper, which is placed as the first column of AP in each table and marked in italics.
@@
In the ablation experiments, ResNet-34 [50] is used as the backbone. The disparity dimensions of multi-scale cost volumes are set to 24, 48, and 96, respectively, from highest-resolution to lowest-resolution. The experiments on KITTI test set use ResNet-50 as the backbone, the disparity dimensions are 48, 48, and 96, respectively. The rest hyperparameters remain unchanged.
@@ IV-B1 Design of positional encoding
Qualitatively, we also illustrate the characteristics of our DAPE in Fig. 4. Given an input left image, we sample a foreground pixel (yellow circle in Fig. 4-(a)) and a background pixel (blue circle). The disparity of the foreground pixel is approximately $12$, thus the DAPE heatmap at $d=12$ is visualized as (b). For the foreground and background pixels, their positional encodings are respectively dot-producted by the encodings at all pixels to emulate the similarity-based attention mechanism. The resultant foreground and background positional encoding similarity heatmaps are visualized as (c) and (d), respectively. To see the attended areas more clearly, we respectively mask the input left image with those heatmaps in (e) and (f). Results show that our DAPE can encode 3D information into the stereo features and attends well on the foreground objects, since the disparities on the surface of an object are of similar distributions.

### v4

@@ Abstract
Transformers have shown promising progress in various visual object detection tasks, including monocular 2D/3D detection and surround-view 3D detection. More importantly, the attention mechanism in the Transformer model and the 3D information extraction in binocular stereo are both similarity-based. However, directly applying existing Transformer-based detectors to binocular stereo 3D object detection leads to slow convergence and significant precision drops. We argue that a key cause of that defect is that existing Transformers ignore the binocular-stereo-specific image correspondence information. In this paper, we explore the model design of Transformers in binocular 3D object detection, focusing particularly on extracting and encoding task-specific image correspondence information. To achieve this goal, we present TS3D, a Transformer-based Stereo-aware 3D object detector. In the TS3D, a Disparity-Aware Positional Encoding (DAPE) module is proposed to embed the image correspondence information into stereo features. The correspondence is encoded as normalized sub-pixel-level disparity and is used in conjunction with sinusoidal 2D positional encoding to provide the 3D location information of the scene. To enrich multi-scale stereo features, we propose a Stereo Preserving Feature Pyramid Network (SPFPN). The SPFPN is designed to preserve the correspondence information while fusing intra-scale and aggregating cross-scale stereo features. Our proposed TS3D achieves a 41.29% Moderate Car detection average precision on the KITTI test set and takes 88 ms to detect objects from each binocular image pair. It is competitive with advanced counterparts in terms of both precision and inference speed.
@@ I Introduction
Stereo vision systems can perceive 3D scene for autonomous vehicles and intelligent transportation [1]. In binocular stereo based 3D perception, image correspondence information plays an essential role [2, 3, 4]. Image correspondence information, in the form of disparity or depth supervision, is widely used in convolutional neural network (CNN) based binocular 3D object detectors to prevent the detector from converging to a monocular local minimum [5, 6].
Vision Transformers [7], which are equipped with similarity-based attention mechanisms and long-range modeling ability, seem compatible with binocular detection at first glance. On the one hand, the attention mechanism in Transformers and the critical 3D information extraction in binocular stereo are both similarity-based. On the other hand, the Transformer’s inherent ability of long-range modeling suites for capturing the wide range of disparities (e.g., a range of $(0,192]$ in KITTI [8]). In addition, Transformers have gained promising progress in 2D object detection [9, 10, 11], surround-view 3D object detection [12, 13, 14], and depth estimation [15]. Therefore, we focus on adopting the Transformer model to binocular 3D detection in this paper, which can hopefully benefit future studies by enabling binocular detection to interact with Transformer-based language models. However, for binocular 3D object detection in driving scenes, to the best of our knowledge, there have not been any public Transformer-based detectors in the literature.
@@
An intuitive way to introduce Transformer models into binocular 3D object detection is to adopt the above surround-view Transformer-based 3D detectors to the binocular scene (e.g., KITTI [8]). To achieve that, we train binocular revisions of existing surround-view 3D object detectors, which are named with Binocular suffixes. Their validation APs after training 320 epochs are enumerated in Fig. 1b, and the validation AP curve of DETR3D-Binocular during training is shown in Fig. 1a as blue circles. It can be observed that existing Transformer detectors converge to a poor 3D detection performance, which can not be compared to advanced CNN-based binocular 3D object detectors nor their surround-view detection performances [16].
We argue that the convergence problem is caused by the lack of image correspondence information in surround-view 3D object detection Transformers such as DETR3D, as well as in existing monocular-based 3D object detection Transformers. It is a fact that monocular images do not involve cross-view correspondence [17, 18]; and the overlapping area across images in a surround-view system is small [16]. Concerning the binocular vision system, however, the overlapping areas between left and right images occupy a large proportion of the pixels. The latent image correspondence information in the large overlap areas is crucial for 3D scene perception. Therefore, extracting and exploiting the image correspondence information in a Transformer-based model is a task-specific challenge of binocular 3D object detection.
@@
Decoder-wisely, a deformable Transformer decoder [21] is introduced to further aggregate matching costs and decode multi-scale stereo features. In the decoder, the DAPE is designed to explicitly encode the sub-pixel-level correspondence information into the stereo features. DAPE composites normalized disparities and sinusoidal 2D positional encoding [21], thus simultaneously providing position information in both 2D image space and 3D scene. In addition to the DAPE, a Non-Parametric Anchor-Based Object Query (NPAQuery) scheme is introduced. Compared with widely-used parametric queries, NPAQuery does not introduce new learnable parameters as query embeddings, instead, it reuses the stereo features as a uniformly distributed query embedding in the 2D image space.
Experiments on the KITTI dataset show that TS3D is an effective Transformer model for binocular 3D object detection, and it is competitive with existing CNN-based counterparts in both terms of accuracy and efficiency. Consequently, our TS3D can encode the image correspondence information in the Transformer model, thus alleviating the above convergence problem. The validation AP curve of our TS3D is plotted in Fig. 1a as orange triangles, and its validation results are listed in Fig. 1b. It is shown that after preserving and encoding the image correspondence information by the proposed SPFPN and DAPE respectively, the performance gap is significantly diminished. To the best of our knowledge, the proposed TS3D is the first public Transformer-based binocular 3D object detector for driving scenes in the literature.
@@
    We present the Transformer-based Stereo-aware 3D object detector (TS3D), which is the first public Transformer-based binocular detector in the literature and can preserve and encode the image correspondence information in binocular images.
@@ II Related work
Existing binocular 3D object detectors are based on convolutional neural networks (CNN) and are categorized in this section. Representative Transformer-based 3D detectors for monocular images, LiDAR point clouds, and surround-view images are briefly introduced, and readers are encouraged to refer to [22, 23] for more thorough reviews on Transformers-based 3D object detection. Existing feature pyramid networks are also reviewed, which relates to our SPFPN.
Existing binocular 3D object detectors can be divided into two categories based on their sources of image correspondence information: detectors with and without LiDAR-based image correspondence supervision. The former [24, 6, 25], shortened as Stereo-with-LiDAR, generates ground-truth image correspondence information based on projected LiDAR point clouds, resulting in accurate yet sparse supervision; whereas the latter [26, 27, 28], shortened as Stereo-without-LiDAR, does not rely on LiDAR point clouds to generate ground-truth image correspondence information. However, the pseudo ground-truths obtained by external stereo matching networks can provide inaccurate supervision in those methods [5, 29]. Some detectors in this category do not use explicit image correspondence supervision [30, 31]. Although Stereo-with-LiDAR detectors achieve higher detection precision, developing Stereo-without-LiDAR detectors is still essential as LiDAR devices may not always be affordable or available, such as in stereo endoscopes and traffic cameras.
@@
The Transformer-based 2D object detector DETR [9] re-formulates the 2D object detection as a collection prediction task and removes the anchor generation and Non-Maximum Suppression (NMS) processes, resulting in an end-to-end detector. DeformableDETR [21] introduces a deformable attention module, which merely applies the multi-head attention mechanism at several learnable key locations around the reference point. It significantly reduces the computational cost of the DETR decoder and converges faster. The Transformer-based unary 3D object detector MonoDTR [17] utilizes auxiliary supervision to implicitly learn depth-aware features and a depth-aware Transformer to mine context and depth information from global features. MonoDETR [43] takes 3D object candidates as queries and introduces an attention-based depth encoder to encode depth embeddings. A depth cross-attention module is proposed to decode inter-query and query-scene interactions. The Transformer-based surround-view 3D object detector DETR3D [12] uses a CNN to extract unary features of each view and a DETR decoder to predict objects. PETR [13] extends DETR3D by generating a 3D coordinate grid in the camera frustum space and fusing the 3D coordinates with the 2D features. BEVFormer [14] generates object queries on the bird’s eye view (BEV) and uses both spatial and temporal information. A spatial cross-attention module is proposed to extract BEV features from each view and a temporal self-attention module is designed to fuse historical BEV features. In addition to the obvious differences in input modalities comparing our TS3D with the above Transformer-based detectors, the image correspondence information is preserved and encoded into enriched stereo feature pyramids by two novel components, namely SPFPN and DAPE.
@@
Feature Pyramid Network (FPN) [44] exploits the multi-level nature of CNNs to generate multi-scale features, so that the features of each level focus more on objects of the specific scale. FPN improves the detection performance of multi-scale 2D object detection. RetinaNet [45] makes use of more levels of the features than that used in FPN, resulting in higher detection precision at a lower computational cost. Original FPN fuses features from lower-resolution to higher-resolution (i.e., top-down), and PAFPN [46] adds an aggregation network from higher-resolution to lower-resolution features (i.e., bottom-up). ASFF [47] adaptively fuses multi-scale features and spatially filters information that conflicts across scales. BiFPN [48] designs a simple and efficient bidirectional FPN, in which a basic structure consisting of top-down and bottom-up paths is repeated to capture enriched multi-scale features. RFP [49] recursively reuses the FPN structure and introduces the Atrous Spatial Pyramid Pooling (ASPP) module as a connection module between two FPNs, which improves the performance of 2D object detection. The above feature pyramids are carefully designed for unary 2D object detection and do not consider the image correspondence nature of binocular images. As a result, they can be applied to generate unary feature pyramids in binocular 3D object detection, nevertheless, directly applying them in stereo features will destroy the latent image correspondence information and affect the accuracy of 3D object detection.
@@ III-A Overall architecture of TS3D
The above unary features are then fed into Stereo Preserving Feature Pyramid Network (SPFPN). Since Transformer-based multi-scale stereo matching is time and memory consuming [15], we utilize multi-scale correlation-based cost volumes [51] for matching cost computation. Both primary and enhanced unary feature pyramids are exploited to construct correlation-based cost volumes, resulting in primary and enhanced cost volume pyramids (shown as orange boxes in the second column of Fig. 2). Constructed cost volume pyramids are then referred to as stereo feature pyramids in the SPFPN. Subsequently, primary and enhanced stereo features from each level of the two pyramids are merged, in which the correspondence information is kept intact. The resultant stereo-preserving feature pyramid is then cross-scale aggregated in a bottom-up manner. Finally, enriched and aggregated multi-scale stereo features are obtained for subsequent disparity estimation and Transformer decoder.
@@
The parametric attention module is used for further cost aggregation and multi-scale detection feature extraction. The Transformer decoder consists of Multi-Head Self-Attention layers [9] (shortened as MHSA in Fig. 2), Multi-Scale Deformable Cross-Attention layers [21] (MSDeformCA), and Feed-Forward Networks (FFNs). The last feature tensor for disparity regression is termed disparity logits and is used as the input of our DAPE. The generated positional encoding of the 3D scene is summed with the query of each layer in the Transformer decoder. The lowest-resolution features in the above feature pyramid are convolved to generate embeddings of the NPAQuery. In addition to the DAPE and NPAQuery, the multi-scale stereo features extracted by SPFPN are exploited as keys and values of the MSDeformCA layers. The above processes including MHSA, MSDeformCA, and FFN are collectively referred to as a decoder layer. The Transformer decoder is a cascade of $N_{\text{dec}}$ decoding layers, in which the object queries are refined step-by-step. During training, 3D objects are classified and regressed under ground-truth supervision after each decoder layer; During inference, only predictions from the last layer are taken as the final 3D object detection results.
@@ III-B Stereo Preserving Feature Pyramid Network
(a) FPN
(b) BiFPN
@@ III-C1 Disparity-Aware Positional Encoding (DAPE)
It can be seen from Eqs. 5 and 6 that the DAPE not only encodes 2D information in the 2D image space but also explicitly encodes the correspondence information of binocular images. It is suitable for locating the projection of 3D object center points and bounding boxes in 2D images; The encoded image correspondence information is guided by stereo features, and thus is suitable for regressing depths and sizes of 3D objects in 3D space.
@@ IV-A Experimental setup
The experiments are conducted on the KITTI dataset [8]. In the ablation experiments, the data split is consistent with that of [55] and the evaluation index is obtained on its validation subset. We report all difficulty levels (Easy, Moderate, Hard) of 3D Car detection AP (Average Precision) as benchmarking indicators. In the experiments in Sec. IV-C, the entire training set is used during training, and the prediction results are calculated by the KITTI server. Detection APs of three difficulty levels of 3D Car and Pedestrian detection tasks are reported on the KITTI test set. Among them, Moderate 3D Car detection AP is the main evaluation index in this paper, which is placed as the first column of AP in each table and marked in italics.
@@
In the ablation experiments, ResNet-34 [50] is used as the backbone. The disparity dimensions of multi-scale cost volumes are set to 24, 48, and 96, respectively, from highest to lowest resolution. The experiments on the KITTI test set use ResNet-50 as the backbone, the disparity dimensions are 48, 48, and 96, respectively. The rest hyperparameters remain unchanged.
@@ IV-B1 Design of positional encoding
Qualitatively, we also illustrate the characteristics of our DAPE in Fig. 4. Given an input left image, we sample a foreground pixel (yellow circle in Fig. 4a) and a background pixel (blue circle). The disparity of the foreground pixel is approximately $12$, thus the DAPE heatmap at $d=12$ is visualized as (b). For the foreground and background pixels, their positional encodings are respectively dot-producted by the encodings at all pixels to emulate the similarity-based attention mechanism. The resultant foreground and background positional encoding similarity heatmaps are visualized as (c) and (d), respectively. To see the attended areas more clearly, we respectively mask the input left image with those heatmaps in (e) and (f). Results show that our DAPE can encode 3D information into the stereo features and attends well on the foreground objects, since the disparities on the surface of an object are of similar distributions.